Work Orders in Molecular Biology Building

Showing all active work orders and all inactive work orders closed in the past 30 days.
WO # Requested by Description Status
155439 Peter Lelonek Condensate Return Pump CRP1 is not working currently. There is water/condensate all over in the room. Completed
156601 Lisa Ristvedt Molecular Biology Building-Room 0121-Remodel Open
156602 Lisa Ristvedt Miscellaneous Open
175503 Lisa Ristvedt Document Reviews Open
189594 Lisa Ristvedt Telecommunications Open
189595 Lisa Ristvedt Moving Open
190519 Peter Lelonek west passenger elevator is stuck on first floor Completed
197249 Peter Lelonek There is low water pressure at the drinking fountain in this area. Users believe it started just after the whole building water shutdown a few weeks back. On Hold
201014 FPM Service Desk FY23 MOL-BIO - General Cleaning Open
202557 FPM Service Desk FY23 MOL-BIO - Supplies Open
206825 Peter Lelonek The passenger elevator is no longer operating. Materials-Awaiting
208748 Peter Lelonek FY23 MOL BIO - SLA for additional custodial services in rooms 0112 and 0114. Every 2 weeks to sweep, mop and empty trash. Stand Renew
208904 Peter Lelonek 0112A MBB - Reconfigure plumbing vent pipe to the south above ceiling so filter box door can be accessed to change filter for AHU in room 0112. Open
210152 Peter Lelonek 1182 MBB - Consult with interior designer on best way to arrange office space for 3 workers and possibly purchase some new furniture. Open
213742 Melissa Warg FDC FEES - DO NOT CHARGE TIME Open
215574 Peter Lelonek The closer on this walk in cold room is not working so the door does not close on its own. On Hold
218846 Peter Lelonek C1100 MBB - Provide cost estimate to replace an existing directional sign in this hallway with a larger one with additional information. Open
218940 Melissa Warg FDC MGMT FEES - DO NOT CHARGE TIME Open
219005 Kasandra Province Requesting (1) Rec. Lamp Barrel 4' Large and (1) Packing Tape 2"x55yd for Molecular Biology Assigned
222915 Joseph Stoberl FPM Shops to order, create, install wayfinding sign. Open
230712 Lisa Ristvedt Molecular Biology Building-Elevator Repairs Open
231151 Peter Lelonek C4010 MBB - The drinking fountain in this hallway no longer works correctly. People have asked if it can be replaced with one that has a bottle filler on it. On Hold
232307 Nicholas Smith Hire contractor to troubleshoot freight elevator door problems. Gate door opens and closes like normal but the landing doors only open about 14 inches on both front and rear doors. Completed
233216 Shane CLOUSER Check DI water pump PWP8 tripped breaker and does not sound good. Materials-Awaiting
238105 Justin Abbott Repair ceiling - damage caused by leak Materials-Awaiting
238294 Peter Lelonek Connect waste piping from drain on new sink Completed
246049 William Webster Replace failed piping domestic water system. Open
246374 William Webster Purchase materials for domestic water service. Open
246691 Kasandra Province Central Receiving delivered a fume hood and cabinets to Molecular Biology from Carver Co Lab on 12/15/22. This ticket should be used for that delivery. Assigned
246692 Kasandra Province Contractor is done installing jack box and conduit for new telecom jack. Requesting service for ISU telecom staff to install the needed wiring. Assigned
246786 William Webster Purchase and install 2 new floats and control panel for condensate receiver tank. Open
246819 Janet Schneider end caps on the following rows are missing either screws, spacers, or both. B north side, C both ends, H north end, G north end, F north end Assigned
252329 Peter Lelonek 3131 MBB - Remodel room with the following items: 1. Add 32" framed opening between adjacent room 3133. Frame to be setup for future door. 2. Remove existing solid ceiling and soffit and replace with new lay in grid ceiling. 3. Replace existing lights with new LEDs 4. Replace existing light switch with occupancy switch 5. Add shelving above refrigerator. Shelving provided by department. 6. Remove existing steam and condensate piping in corner of room. 7. Add new power along west wall for equipment. Add 4 separate 120 volt, 20 amp circuits. 8. Add new supply and return grilles Open
252371 Melissa Warg FDC MGMT FEES - DO NOT CHARGE TIME Open
252841 Brady Streit Contractor - Remove and replace corroded domestic piping on all floor. We need a specification and cost for replacement of galvanized piping for the three risers and the mechanical room. Pricing can be per system or total. Should include costs; to support outage (Facilities Services labor). Brian Grove has been working with us on this project. Completed
254630 Peter Lelonek 3182 MBB - The hot water at the faucet just inside the door to this lab does not flow very fast. On Hold
255231 Kasandra Province Telecommunications - Please add telecommunication jacks and cables to Room 0121. There is an empty telecommunications box above the new door, along the South side of it, close to the ceiling line. Open
258865 Kasandra Province Please provide fire extinguisher for newly renovated Molecular Biology Room 0121 Open
258883 Lyle Bishop East penthouse has a detached hanger. 1 blocked sprinkler head in M2134-1. 1 Taped head in north west corner of basement riser room. Open
259360 Christopher Grosz need to thaw and insulate refrigerant piping on cold room condensing unit attn chris grosz On Hold
259500 Nicholas Smith Hire Schumacher elevator to troubleshoot issues with west passenger elevator. possible drive issue. ATTN Tom Webster Assigned
265449 Abigail Zumpano Contracts Open
265450 Abigail Zumpano Fire Alarm Support On Hold
265824 Brandon Thompson From Inspection 5979 - Fume hood E033462 has a low face velocity of 44, Please service to an acceptable range (70-150 fpm). Completed
265825 Brandon Thompson From Inspection 5980 - Fume hood E033459 has a low face velocity of 35, Please service to an acceptable range (70-150 fpm). Completed
266968 Nicholas Smith replace anti-spin cable on top of elevator per inspector ATTN Nick Smith Assigned
267135 Brandon Thompson From Inspection 6093 - Fume hood E033399 has a low face velocity of 43, Please service to an acceptable range (70-150 fpm). Completed
267137 Brandon Thompson From Inspection 6095 - Fume hood E033398 has a low face velocity of 35, Please service to an acceptable range (70-150 fpm). Completed
267138 Brandon Thompson From Inspection 6095 - Fume hood E033398 has a low face velocity of 35, Please service to an acceptable range (70-150 fpm). Completed
267172 Brandon Thompson From Inspection 6095 - Fume hood E033398 has a low face velocity of 35, Please service to an acceptable range (70-150 fpm). Completed
267174 Brandon Thompson From Inspection 6096 - Fume hood E033458 has a low face velocity of 34, Please service to an acceptable range (70-150 fpm). Completed
267246 Peter Lelonek 2288 MBB - Replace all burnt out lamps in this lab. Assigned
267347 Brandon Thompson From Inspection 6168 - Fume hood E033434 has a low face velocity of 42. Please service to an acceptable range (70-150 fpm). Completed
268205 Brandon Thompson From Inspection 6204 - Fume hood E033004 has a low face velocity of 39. Please service to an acceptable range (70-150 fpm). Completed
268361 Brandon Thompson From Inspection 6233 - Fume hood E033442 tested fine but the sash is extremely difficult to raise and lower and needs adjusting. Nothing seems to be wrong with the cables, but the left side of the lower part of the sash does scrape against the side of the hood when trying to move it. On Hold
268368 Lawrence Woolard Lights nearest stall need to be replaced Assigned
268434 Peter Lelonek The freight elevator is not coming to a floor when called. the elevator works fine if you are on it but you must go to the floor where it is sitting to get on it. Assigned
268461 Peter Lelonek 0124 MBB - Replace all dim and burnt out lamps. Lamp color and brightness needs to be consistent in the room due to the science being done. Assigned
268462 Peter Lelonek 1160A MBB- The DI faucet in this room is not working. Assigned
268524 Rose Perksen Water leak coming rom ceiling Completed
272317 Peter Lelonek 3137 MBB - Add emergency power 208V, 20 amp, 1 phase circuit to west side of room in open wiremold area for -80 freezer. I am hoping you can reroute wire from an existing 208V circuit already in the room but feeding from an adjacent panel. The old circuit is 37/39 in panel 31C and the new circuit needs to feed from panel 31E. The outlet for this old circuit in located on the south wall of room 3137 at the east end of the wiremold. Assigned
272402 Donald Gregory DI water feeding pure water pump 8 is leaking Materials-Awaiting
272465 Peter Lelonek 0114 MBB - This lab room is over cooling. Assigned
272473 Michael Olson CP941665-Utility Engineering-Molecular Biology room 0121 Open
272870 Peter Lelonek 4181 MBB - This cold room has warmed up again. Completed
273059 Todd Kelley replace VFD Attn Todd Kelley #197120 closed and processed Open
273213 Kasandra Province Repair fume hood exhaust flexible duct at tear Open
273274 Peter Lelonek 3232, 3238, 3242 MBB - replace all dim or burnt out lamps in these 3 labs Assigned
273473 Lawrence Woolard Lights nearest to the toilet stall need to be replaced Assigned
273676 Peter Lelonek AHU4 is not opening the cooling water valve as needed to maintain discharge air temperature. I think the valve is hung up in METASYS and needs to be triggered. This same thing happened last year. Can a point be setup in METASYS so I can trigger this process in the future without needing to turn in a request? Completed
273683 Michael Fields Environmental Chambers at Molecular Biology Building - Service work to be completed on the following units: Room1243 / RSE10 - middle evaporator fan not running Room 2243 / RSE13 - door closure adjustment required Room 2235 / RSE2 - evaporator fan not running Room 2293 / RSE12 - door closure adjustment required Room 2181A / RSE17 - door closure adjustment required Room 3241 / RSE20 - door closure adjustment required Room 3191 / RSE23 - door seal replacement required Completed
273788 Aaron Crossett Bulbs are out in both fixtures Assigned
273815 Peter Lelonek 1210 MBB - Replace all lamps that are dim or burnt out in this admin area. Assigned
279138 Peter Lelonek 3282 and 3281 MBB - Coordinate cleaning and waxing of lab floors in these rooms with Olga Zabotina in office 3212 Assigned
279210 Peter Lelonek Modify occupied schedule for operation of AHU4 to shutoff at 8pm Mon thru Friday and 6pm Sat/Sun.Start times can remain as they are. Assigned
279265 Peter Lelonek The freight elevator still has issues being called to some floors. It works once in it but doesn't always come to the floor when called. Assigned
279356 Peter Lelonek Fire panel is showing 3 troubles. Assigned
279588 Abigail Zumpano Remove and cap old steam pipe. See Eric Ohrt or Pete Lelonek for details. Attn: Kemble or Solberg Assigned
279801 Peter Lelonek Greenhouse room is overheating, and the vent are not opening. Assigned
279858 Kevin Vols Vesda Trouble Open Circuit Assigned
280128 Timothy Kelley LVL 1 N.W. Lobby Smoke Detector IO51 Smoke detector Fire Alarm- Cleaned head and reset Fire panel Call back Att. Tim Kelley Assigned
280199 Peter Lelonek 2360 MBB - this greenhouse is overheating because the cooling dampers don't open. Assigned
280300 Peter Lelonek The cooling valve on AHU-1 is not opening up when cooling is needed. Seems to be hung up in METASYS controls. It has been overridden until it can be corrected. Assigned
280548 Lyle Bishop Hire Interstate Power Systems to perform Oil & filter change on emergency generator. ATTN: Lyle Bishop. Assigned
280590 John Breon Contractor to give estimate to repair roof leak Open
280627 Peter Lelonek 3264 MBB - Paint academic office using outside contractor. Color selected by user. Open
280648 Melissa Warg FDC MGMT FEES - DO NOT CHARGE TIME Open
280710 Peter Lelonek Make and install approximately 17 new or update some existing gray plastic signs with white lettering to improve way finding of Biotech Facilities on ground and first floor. Sign locations and wording is already determined. I need a cost estimate before proceeding. Open
280911 Melissa Warg FDC MGMT FEES - DO NOT CHARGE TIME Open
280992 Peter Lelonek 2191 MBB - There is an outlet in the middle of the south wall of this room that stopped working. There is not any breakers tripped that we can find for this room. On Hold
281083 Eric Zierke Improvements for outside air intakes Assigned
281205 Peter Lelonek There is a steam trap that has a hole in it that is leaking steam out. It is located a little bit to the south of the center of the mechanical room and above the GA11 air compressor. We have a shutdown of the steam system that this trap is on scheduled for tomorrow morning. It would be good if this could be replaced at the same time as that shutdown. Assigned
281260 Michael Fields Moly Bio Bld - Room 0240 AHU 1 - Chilled water control valve leaking water onto the floor. (constant dripping from valve stem area) Assigned
281432 Peter Lelonek 1132 MBB - There is a drip from the steam line serving the autoclave in this room. The drip is located before the isolation valve at the ceiling on the left side of the autoclave. Assigned
281433 Peter Lelonek The lights on the dock area at the east side of the building are on all of the time. The light in the same area above the room up garage door does not seem to work and may need a new lamp. Assigned
281599 Abigail Zumpano Contracts Open